# SuperQuadricOcc: Multi-Layer Gaussian Approximation of Superquadrics for Real-Time Self-Supervised Occupancy Estimation

### v1

@@ IV-B Implementation
For a fair comparison with GaussianFlowOcc [boeder2025gaussianflowocc], we train our model following their experimental configuration. We use a ResNet-50 backbone with an input image resolution of $256\times 704$ and adopt their Gaussian Transformer network comprising three blocks. Induced attention is implemented with 500 inducing points. Our approach employs 1600 superquadric primitives, compared to the 10,000 Gaussians used in GaussianFlowOcc. In the superquadric-to-Gaussian module, we use an 80-faced icosahedron with scale values: $K=[0.5,0.6,0.75,0.9,1.05,1.2,1.6,2.0,2.5]$, forming an $S_{K=9}$ scaled family of superquadrics. This results in 720 Gaussians per superquadric for rendering. All models are trained for 18 epochs on 4 NVIDIA A100 GPUs with a batch size of 6.
@@ IV-C Main Results
In this section, we present the main results of our model compared to GaussianFlowOcc on the Occ3D-nuScenes [tian2023occ3d] and OpenOcc [sima2023scene] datasets. Table I reports various results on Occ3D, where our model surpasses the Gaussian-based GaussianFlowOcc, achieving a 5.9% increase in mIoU while utilizing 84% fewer primitives. When GaussianFlowOcc is trained with an equivalent number of primitives (1600), performance drops by 16.7% in mIoU compared to its 10,000-Gaussian configuration. This highlights the spatial expressiveness of superquadrics due to their diverse shape parameterization. However, we observe a reduction in IoU relative to GaussianFlowOcc, likely caused by a slight overdensification resulting from a mismatch between the Gaussian loss representation and the superquadric evaluation representation. Furthermore, SuperQuadricOcc outperforms GaussianFlowOcc on dynamic object detection, particularly for vulnerable road users, showing a +3.84 IoU improvement for the bicyclist class. This gain stems from the irregular shapes of such objects, which are better captured by the flexible geometry of superquadrics. Although GaussTR [jiang2025gausstr] achieves higher mIoU, this is attributed to its use of foundation models during inference, which significantly degrades its inference speed, limiting its applicability for real-time systems.
@@
TABLE II: RayIoU evaluated on the Occ3D-nuScenes [tian2023occ3d] dataset: GaussianFlowOcc* denotes the model trained with 1,600 Gaussians. The best-performing model for each evaluation metric is highlighted in bold.
Model RayIoU RayIoU@1 RayIoU@2 RayIoU@4 GaussianFlowOcc [boeder2025gaussianflowocc] 10.38 6.49 10.01 14.65 GaussianFlowOcc* [boeder2025gaussianflowocc] 9.69 5.79 9.40 13.88 SuperQuadricOcc (Ours) 10.93 6.62 10.70 15.49
TABLE III: RayIoU evaluated on the OpenOccv2 [tian2023occ3d] dataset: GaussianFlowOcc* denotes the model trained with 1,600 Gaussians. The best-performing model for each evaluation metric is highlighted in bold.
Model RayIoU RayIoU@1 RayIoU@2 RayIoU@4 GaussianFlowOcc [boeder2025gaussianflowocc] 11.86 7.78 11.61 16.19 GaussianFlowOcc* [boeder2025gaussianflowocc] 10.71 6.64 10.48 15.01 SuperQuadricOcc (Ours) 12.46 7.78 12.29 17.30
@@ IV-E Visualizations
Figure 4: Visualization of Primitives and Voxel Estimation: Comparison between SuperQuadricOcc and the 10,000- and 1600-GaussianFlowOcc models on the Occ3D-nuScenes dataset [tian2023occ3d] for sample tokens edde57e6dfeb416e936d9056824b8253 and 135bf33890ba4ca2984a931444923eda.
@@ A-A Semantic Occupancy
Furthermore, we provide videos comparing the occupancy predictions of SuperQuadricOcc with those of the 1600- and 10,000-Gaussian GaussianFlowOcc models [boeder2025gaussianflowocc], alongside ground truth, for three scenes: scene-0016, scene-0092, and scene-0276. Consistent with earlier analysis, SuperQuadricOcc produces superior predictions due to its diverse shape set, particularly when compared with the 1600-Gaussian GaussianFlowOcc.
@@ A-B Depth and Semantic Renders
Figure 6: Visualization of Rendered Semantics and Depth: Comparison of semantic and depth renderings produced by SuperQuadricOcc and 10,000-GaussianFlowOcc on the Occ3D-nuScenes dataset [tian2023occ3d] for sample token 05cf2a68a2cf4ff1838034f8ecae0c82.

### v3

@@ IV-B Implementation
For a fair comparison with GaussianFlowOcc [4], we train our model following their experimental configuration. We use a ResNet-50 backbone with an input image resolution of $256\times 704$ and adopt their Gaussian Transformer network comprising three blocks. Induced attention is implemented with 500 inducing points. Our approach employs 1600 superquadric primitives, compared to the 10,000 Gaussians used in GaussianFlowOcc. In the superquadric-to-Gaussian module, we use an 80-faced icosahedron with scale values: $K=[0.5,0.6,0.75,0.9,1.05,1.2,1.6,2.0,2.5]$, forming an $S_{K=9}$ scaled family of superquadrics. This results in 720 Gaussians per superquadric for rendering. All models are trained for 18 epochs on 4 NVIDIA A100 GPUs with a batch size of 6.
@@ IV-C Main Results
In this section, we present the main results of our model compared to GaussianFlowOcc on the Occ3D-nuScenes [40] and OpenOcc [36] datasets. Table I reports various results on Occ3D, where our model surpasses the Gaussian-based GaussianFlowOcc, achieving a 5.9% increase in mIoU while utilizing 84% fewer primitives. When GaussianFlowOcc is trained with an equivalent number of primitives (1600), performance drops by 16.7% in mIoU compared to its 10,000-Gaussian configuration. This highlights the spatial expressiveness of superquadrics due to their diverse shape parameterization. However, we observe a reduction in IoU relative to GaussianFlowOcc, likely caused by a slight overdensification resulting from a mismatch between the Gaussian loss representation and the superquadric evaluation representation. Furthermore, SuperQuadricOcc outperforms GaussianFlowOcc on dynamic object detection, particularly for vulnerable road users, showing a +3.84 IoU improvement for the bicyclist class. This gain stems from the irregular shapes of such objects, which are better captured by the flexible geometry of superquadrics. Although GaussTR [15] achieves higher mIoU, this is attributed to its use of foundation models during inference, which significantly degrades its inference speed, limiting its applicability for real-time systems.
@@
TABLE II: RayIoU evaluated on the Occ3D-nuScenes [40] dataset: GaussianFlowOcc* denotes the model trained with 1,600 Gaussians. The best-performing model for each evaluation metric is highlighted in bold.
Model RayIoU RayIoU@1 RayIoU@2 RayIoU@4 GaussianFlowOcc [4] 10.38 6.49 10.01 14.65 GaussianFlowOcc* [4] 9.69 5.79 9.40 13.88 SuperQuadricOcc (Ours) 10.93 6.62 10.70 15.49
TABLE III: RayIoU evaluated on the OpenOccv2 [36] dataset: GaussianFlowOcc* denotes the model trained with 1,600 Gaussians. The best-performing model for each evaluation metric is highlighted in bold.
Model RayIoU RayIoU@1 RayIoU@2 RayIoU@4 GaussianFlowOcc [4] 11.86 7.78 11.61 16.19 GaussianFlowOcc* [4] 10.71 6.64 10.48 15.01 SuperQuadricOcc (Ours) 12.46 7.78 12.29 17.30
@@ IV-E Visualizations
Figure 4: Visualization of Primitives and Voxel Estimation: Comparison between SuperQuadricOcc and the 10,000- and 1600-GaussianFlowOcc models on the Occ3D-nuScenes dataset [40] for sample tokens edde57e6dfeb416e936d9056824b8253 and 135bf33890ba4ca2984a931444923eda.
@@ V Conclusion
-A Render Time
In Table IX, we demonstrate minimal memory usage due to the adoption of Gaussians for rendering. Further, render time is 246ms to render six camera views for a given sample. Inference time could be reduced by reducing the number of Gaussian approximation layers.
TABLE IX: Inference Results: We present render time and memory usage average over the validation set.
Memory (GB) Inference Time (ms) 1.35 246.83
@@ A-A Semantic Occupancy
Furthermore, we provide videos comparing the occupancy predictions of SuperQuadricOcc with those of the 1600- and 10,000-Gaussian GaussianFlowOcc models [4], alongside ground truth, for three scenes: scene-0016, scene-0092, and scene-0276. Consistent with earlier analysis, SuperQuadricOcc produces superior predictions due to its diverse shape set, particularly when compared with the 1600-Gaussian GaussianFlowOcc.
@@ A-B Depth and Semantic Renders
Figure 6: Visualization of Rendered Semantics and Depth: Comparison of semantic and depth renderings produced by SuperQuadricOcc and 10,000-GaussianFlowOcc on the Occ3D-nuScenes dataset [40] for sample token 05cf2a68a2cf4ff1838034f8ecae0c82.